### v2

# Restore Text First, Enhance Image Later: Two-Stage Scene Text Image Super-Resolution with Glyph Structure Guidance

Abstract
Current image super-resolution methods show strong performance on natural images but distort text, creating a fundamental trade-off between image quality and textual readability. To address this, we introduce TIGER (Text–Image Guided supEr-Resolution), a novel two-stage framework that breaks this trade-off through a “text-first, image-later” paradigm. TIGER explicitly decouples glyph restoration from image enhancement: it first reconstructs precise text structures and uses them to guide full-image super-resolution. This ensures high fidelity and readability. To support comprehensive training and evaluation, we present the UZ-ST (UltraZoom-Scene Text) dataset, the first Chinese scene text dataset with extreme zoom. Extensive experiments show TIGER achieves state-of-the-art performance, enhancing readability and image quality.
Figure 1: We present TIGER (Text–Image Guided supEr-Resolution), a novel framework for scene text super-resolution. Its ‘text-first, image-later’ paradigm ensures accurate glyph restoration and consistently high overall image fidelity and visual quality.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
Scene Text Image Super-Resolution is a critical problem in computer vision for its vital role in car navigation [DBLP:conf/iclr/LiC25], scene understanding [kil2023prestu, deshmukh2024textual], and document enhancement [souibgui2023text]. It seeks to restore a high-quality super-resolution (SR) image from a degraded low-resolution (LR) input while preserving the text meaning. Unlike general image super-resolution for natural scenes, where plausible texture synthesis or detail hallucination is acceptable and does not alter the semantic meaning of the image, visual texts exhibit extremely low tolerance to structural errors. This is especially evident in Chinese, where a minor stroke distortion or omission may entirely alter the meaning [yu2023chinese, li2023learning].
In recent years, with the rapid development of generative models such as diffusion [ho2020denoising, dhariwal2021diffusion, saharia2022photorealistic, ramesh2022hierarchical, rombach2022high, chang2023muse, zhang2024artbank], image super-resolution has increasingly relied on their strong generative priors to recover missing details from low-resolution (LR) inputs [wu2024seesr, yu2024scaling, lin2024diffbir, wu2025one, DBLP:conf/cvpr/DongFGWZCLZ25, duan2025dit4sr, hu2025text]. Although these approaches can effectively restore fine-grained natural textures (e.g., grass, leaves), they often distort text regions, turning them into gibberish as shown in Fig.1 (b). Because their primary goal is overall image enhancement, they tend to overlook text structure, resulting in a limited ability to preserve fine-grained glyph details. This limitation is amplified for Chinese characters, due to their complex and diverse glyph designs and relatively low saliency in images. Consequently, models tend to under-represent and collapse glyphs into oversimplified, averaged forms, producing overlapping or distorted characters. Some researchers attempt to address this by processing only text regions [zhang2024diffusion, li2023learning]. While these methods improve readability notably, the absence of global background constraints often introduces style inconsistencies and block artifacts between text and background. Ultimately, current approaches face a persistent trade-off between maintaining readability and ensuring high image quality. Our key observation is that these goals need not be mutually exclusive if text and non-text are explicitly handled differently. Text structures can be reconstructed with dedicated mechanisms and then used to guide full-image restoration, ensuring coherent style without artifacts.
Building on this insight, we introduce TIGER (Text–Image Guided supEr-Resolution), a progressive two-stage paradigm for scene text super-resolution built on the principle of “restoring text structure first, enhancing the whole image later.” Unlike methods that rely on a single generative prior, TIGER explicitly separates the treatment of text and non-text regions: a diffusion-based local text refiner focuses on reconstructing fine-grained stroke geometry in text regions, ensuring glyph fidelity and structural consistency. The recovered text structures are then injected as conditional guidance into the subsequent full-image restoration stage, steering global super-resolution to harmonize text and background while suppressing artifacts and preserving overall visual quality.
Table 1: Dataset comparison. TextZoom [wang2020scene] lacks Chinese, CTR [yu2021benchmarking] lacks multi-line text, and Real-CE [ma2023benchmark] has mild degradation. UZ-ST covers all, enabling comprehensive evaluation.
Content TextZoom CTR Real-CE UZ-ST (Ours)
Chinese characters
Multi-line text
Real-World Degradation($\times 4$)
Moreover, existing scene text super-resolution datasets [wang2020scene, yu2021benchmarking, ma2023benchmark] are limited to mild degradation ($\leq\!\times 4$), which fails to reflect real-world long-distance scenarios and restricts models from learning robust low- to high-resolution mappings. To address the issue of data scarcity and enable comprehensive evaluation [wang2020scene, ma2023benchmark], we introduce the UZ-ST (UltraZoom-Sence Text) benchmark dataset, the first scene text dataset that contains extreme zooming mode ($\times 14.29$), providing extra challenging scenarios for the field. Content differences are outlined in Tab. 1. It includes high-quality 5,036 LR–HR pairs captured at multiple focal lengths (ranging from 14mm to 200mm), with 49,675 text lines in total. Each pair comes with detailed annotations, including detection boxes and text transcripts, to support both training and evaluation. The dataset includes diverse scenarios such as shop signs, posters, and documents. It also covers varying lighting conditions, including daylight, indoor lighting, and nighttime, offering a challenging yet practical benchmark for the field. To ensure reliable alignment across extreme focal lengths, we adopt a coarse-to-fine cascade alignment strategy combined with manual filtering: coarse alignment is performed via global geometric transformation, and residual misalignments are corrected through local refinement. This process achieves accurate LR–HR alignment, thereby ensuring both the quality of the dataset and the reliability of our experiments. The main contributions are summarized as follows:
  1. 1.
    We propose TIGER, the first two-stage scene text super-resolution framework with a ‘text-first, image-later’ paradigm that decouples glyph restoration from image enhancement, improving readability and visual quality.
  2. 2.
    We introduce UZ-ST, the first scene text benchmark with extreme zoom ($\times$14.29), offering aligned, richly annotated LR–HR pairs for thorough evaluation under challenging real-world conditions.
  3. 3.
    Extensive experiments on Real-CE and UZ-ST show our method outperforms prior state-of-the-art models, especially in preserving text structure.
2 Related Works
Figure 2: The framework of TIGER, which includes the Text Restoration stage (stage 1) and the Image Enhancement stage (stage 2). Stage 1 recover accurate glyph structures from text regions. Stage 2 uses them to guide full-image restoration for coherent text and background.
Real-World Image Super-Resolution.  Real-world image super-resolution (Real-SR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs degraded under uncontrolled real-world conditions. Early GAN-based methods, such as BSRGAN [zhang2021designing] and Real-ESRGAN [wang2021real], synthesize degradations through random combinations of known distortions and use adversarial training for restoration. While they generate natural-looking images, their instability and insensitivity to fine details limit their ability to recover structural elements, especially text. Recent work introduces diffusion models into Real-SR [yue2023resshift, wang2024sinsr], improving perceptual quality but still struggling with complex structures like scene text. StableSR [wang2024exploiting] and DiffBIR [lin2024diffbir] apply ControlNet [zhang2023adding] to condition generation on LR inputs, while PASD [yang2024pixel] and SeeSR [wu2024seesr] incorporate high-level semantics to enhance fidelity. SUPIR [yu2024scaling] scales training with large image-text pairs and introduces degradation-robust encoders. Other approaches, including OSEDiff [wu2025one] and TSD-SR [dong2025tsd], directly apply the diffusion process on LR images and distill models for one-step sampling. DiT-SR [cheng2025effective], DreamClear [ai2024dreamclear], and DiT4SR [duan2025dit4sr] adopt diffusion transformers (DiT) for Real-SR. Despite enhancing perceptual fidelity, these methods often overlook accurate text structures. TADiSR [hu2025text] addresses Chinese scene text super-resolution by using Kolors as the base model and aggregating cross-attention maps for text structure supervision. However, constrained by the resolution of cross-attention, it struggles to restore small or severely degraded text.
Text Image Super-Resolution. Text image super-resolution (Text-SR) restores textual content from cropped images of isolated words or text lines. Early methods [dong2015boosting] apply general SR architectures such as SRCNN [dong2014learning] to enhance OCR performance. TextSR [wang2019textsr] introduces GANs with a recognition loss, while PlugNet [mou2020plugnet] and STT [chen2021scene] jointly train SR and recognition modules for more discriminative features. TSRN [wang2020scene] releases the TextZoom dataset and employs an edge-aware module to preserve text details, while TATT [ma2022text] uses global attention for irregular layouts. Methods like MTDM [wang2024multi] inpaint text but fail on complex scripts due to primitive architectures. MARCONet [li2023learning] employs StyleGAN priors and a glyph structure codebook for realistic text reconstruction. DiffTSR [zhang2024diffusion] employs latent diffusion to separately denoise text and text-image components. Despite progress in text structure restoration, existing methods lack global background constraints, causing style inconsistencies and block artifacts between text and background.
3 Methodology
3.1 Architecture Overview
Current image super-resolution methods focus on enhancing overall image quality but often fail to accurately preserve glyph structures, leading to distorted text in super-resolved images, as shown in Fig. 1 (b). Conversely, text image super-resolution methods improve text readability but do not retain the global semantic information of the background, causing incoherence between text and non-text regions. To leverage the strengths of both approaches, we introduce the TIGER framework. As illustrated in Fig. 2, the TIGER framework is composed of 2 stages, the Text Restoration stage and the Image Enhancement stage. In the text restoration stage, we extract the text regions from the LR input $x_{L}\in\mathbb{R}^{H\times W\times C}$ and feed them into the glyph structure restoration model to restore the text structure based on the text region of the LR input and the predicted text. We then reassemble the text structures to their original positions to obtain a text mask $\hat{x}_{m}\in\mathbb{R}^{H\times W\times C}$. In the image enhancement stage, the text mask and LR input are then processed by a ControlNet-like network to obtain the enhanced SR output $\hat{x}_{H}\in\mathbb{R}^{H\times W\times C}$. The following sections detail each stage.
3.2 Text Restoration stage
Figure 3: Overview of UZ-ST (UltraZoom-Scene Text). (a) Real-CE LRs show only mild degradation (red box), while UZ-ST LRs exhibit stronger degradation (red box), enabling a more comprehensive evaluation. (b) Coarse-to-fine alignment: images are sorted by focal length, each warped to the next higher-focal neighbor using an estimated homography matrix, then refined to the 200 mm GT.
Text Restoration Pipeline. Existing methods like Hi-SAM [ye2024hi] work well on clean high-resolution text but fail on the incomplete, distorted text in low-resolution images. To address this, we propose a region-level text restoration pipeline shown in Fig. 2 built on [rombach2022high] to recover glyph structures $x_{m}\in\mathbb{R}^{H\times W\times C}$. An OCR detector first localizes text regions $\{x^{0}_{L},...,x^{N-1}_{L}\}$ and extracts their contents $\{y^{0},...,y^{N-1}\}$ as semantic conditions. Each region $\tilde{x}_{L}$ is encoded by VAE [kingma2013auto] into $\tilde{z}_{L}\in\mathbb{R}^{h\times w\times c}$, concatenated with noise $z_{T}\in\mathbb{R}^{h\times w\times 2c}$, and iteratively denoised by a UNet $\epsilon_{\theta}$ into two branches: $z^{RGB}_{t-1}\in\mathbb{R}^{h\times w\times c}$ for appearance and $z^{m}_{t-1}\in\mathbb{R}^{h\times w\times c}$ for structure, which are merged as $z_{t-1}$. The text content $y$ is embedded into $c_{te}$ and fused into $z_{t}$ via cross-attention to guide structure recovery. After $T$ denoising steps, the mask branch output $z^{m}_{0}$ is decoded into $\tilde{x}_{m}$, and all restored regions are assembled into the final text mask $\hat{x}_{m}$. Focusing on real text reconstruction enables the model to capture glyph structures without non-text interference and reduces sensitivity to text saliency.
Training Strategy. The scarcity of real-world segmentation masks for degraded text forces reliance on synthetic data, yet its artificial degradation limits generalization. Annotating real degraded samples, however, is labor-intensive and costly. To marry synthetic precision with real-world degradation, we propose a two-phase training strategy. In Phase 1, both synthetic and real data are included in training. This allows the model to generate text masks from the LR image and capture real-world degradation patterns. However, noisy masks in real data degrade mask quality. To address this, we freeze the RGB out block and the mask out block and train the UNet with only synthetic data to refine the quality of output text masks in Phase 2. The training objective is described as follows:
$\mathcal{L}=\lambda_{td}\mathcal{L}_{td}+\lambda_{Seg}\mathcal{L}_{Seg},$ (1)
where $\mathcal{L}_{td}$, $\mathcal{L}_{Seg}$, $\lambda_{td}$, and $\lambda_{Seg}$ denote the text-control diffusion loss, segmentation-oriented loss, and their corresponding hyperparameters. The text-control diffusion loss is formulated as follows:
$\mathcal{L}_{\epsilon}=\mathbb{E}_{\textbf{{z}}_{0},\tilde{\textbf{{z}}}_{L},\textbf{{c}}_{te},\textbf{{t}},\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}(\textbf{{z}}_{t},\tilde{\textbf{{z}}}_{L},\textbf{{c}}_{te},\textbf{{t}})\|_{2}^{2}\right].$ (2)
To improve the quality of the text mask, we propose a segmentation-oriented loss. Let $\mathcal{\varepsilon}_{t}$ denote the noise predicted by the denoiser network $\epsilon_{\theta}$. Following  [ho2020denoising], $z^{m}_{0}$ can be estimated by combining the time step $t$ with the noisy latent image $z^{m}_{t}$. This estimate is subsequently passed through the VAE decoder to obtain an approximate reconstruction of the original input text mask, denoted as $x^{\prime m}_{0}$. In this way, text mask generation can be supervised at the pixel level. We combine Mean Squared Error (MSE), Focal, and Dice Losses, commonly used in segmentation tasks [ye2024hi], to compare $x^{\prime m}_{0}$ with the original image $x^{m}_{0}$:
$\mathcal{L}_{Seg}=\|x^{\prime m}_{0}-x^{m}_{0}\|^{2}_{2}+\lambda_{Focal}\text{FocalLoss}(x^{\prime m}_{0},x^{m}_{0})$ (3)
$+\lambda_{Dice}\text{DiceLoss}(x^{\prime m}_{0},x^{m}_{0}),$
where $\lambda_{Focal}$ and $\lambda_{Dice}$ are their respective balancing coefficients. This combination forms a novel text-first paradigm that equips the model with both structural accuracy and real-world generalization, laying the foundation for the subsequent full-image super-resolution stage.
3.3 Image Enhancement stage
Image Enhancement pipeline. To effectively take advantage of the generated glyph structure $\hat{x}_{m}$ and enhance the quality of LR $x_{L}$, we adopt a ControlNet [zhang2023adding] network $\epsilon_{\phi}$. As illustrated in Fig. 2, after getting their latent representation $\hat{z}_{m}$ and $z_{L}$, we use the network to denoise the $z_{L}$ only at the specific timestep $t$, using the null-text embedding $c_{Null}$. The output super-resolved latent representation $\hat{z}_{H}$ will be denoted as:
$\hat{z}_{H}=z_{L}-\sigma_{t}\epsilon_{\phi}(z_{L},\hat{z}_{m},t,c_{Null}),$ (4)
where $\sigma_{t}$ is a scalar determined by the predefined diffusion time step. With the text restoration stage, this image enhancement pipeline forms a progressive ‘text-first, image after’ paradigm. Our design injects structure-aware control into the generative process, allowing the network to enhance global image quality while preserving text structures.
Training Strategy. For the quality of image super-resolution, we constrain the reconstruction loss between the predicted high-resolution image $\hat{x}_{H}$ decoded from $\hat{z}_{H}$ by $\varepsilon$ and the ground truth high-resolution image $x_{H}$ using a weighted sum of MSE and LPIPS losses:
$\mathcal{L}_{img}=\lambda_{l2}\|x_{H}-\hat{x}_{H}\|^{2}_{2}+\lambda_{LPIPS}\text{LPIPS}(x_{H},\hat{x}_{H}),$ (5)
where $\lambda_{l2}$ and $\lambda_{LPIPS}$ are balancing coefficients for different loss terms. To enhance glyph control and emphasize the glyph structure, we extract the text boundaries using Sobel operators [roberts1987digital]. This edge loss is expressed as:
$\mathcal{L}_{edge}=\|\text{Sobel}(x_{H})-\text{Sobel}(\hat{x}_{H})\|^{2}_{2}.$ (6)
We combine the two loss terms and use $\lambda_{edge}$ as the balancing coefficient in the complete loss function for stage 2:
$\mathcal{L}=\mathcal{L}_{img}+\lambda_{edge}\mathcal{L}_{edge}.$ (7)
The combination of glyph-aware ControlNet conditioning and edge-constrained training objectives enables our model to preserve stroke integrity while harmonizing text and background appearance.
4 Dataset and Benchmark
Existing datasets like ma2023benchmark offer only subtle degradation, as shown in Fig. 3 (a), making them insufficient for evaluating model robustness. To address the lack of both challenging and publicly available datasets tailored for scene text image super-resolution, especially for Chinese text, we introduce UZ-ST, a challenging real-world benchmark. It is collected using ViVO X200 Ultra equipped with four fixed focal lengths (14 mm, 35 mm, 85 mm, 200 mm), enabling image pairs with extreme $\times 14.29$ zoom. It includes diverse scenes—street views, book covers, advertisements, menus, and posters—captured under varied lighting, with all text lines manually annotated. However, such extreme zoom introduces severe misalignment that breaks pixel-wise optimization methods like [cai2019toward], originally designed for moderate $\times 2$ to $\times 4$ zoom. To address this, we design a Cascade Coarse-to-fine Alignment pipeline (see Fig. 3 (b)): images are first sorted by focal length, and each low-quality image is sequentially aligned to its next higher-focal neighbor using optimization-based registration or feature-based extractors [lowe1999object, ripe2025], then refined against the 200 mm ground truth. Finally, we manually filter or realign problematic cases to ensure precise alignment.
Table 2: Evaluation results on image quality and text accuracy. Numbers in bold indicate the best performance. TIGER performs best.
Methods Real-CE UZ-ST
[…]
Real-ESRGAN 22.30 0.787 0.239 0.188 53.60 56.0% 23.99 0.790 0.248 0.194 30.60 37.1%
HAT 23.61 0.830 0.214 0.176 51.16 56.6% 25.17 0.815 0.249 0.198 30.12 37.9%
MARCONet 21.89 0.785 0.238 0.150 52.97 55.0% 22.13 0.768 0.306 0.205 34.28 33.4%
SeeSR 23.59 0.822 0.195 0.169 43.75 37.4% 23.64 0.788 0.219 0.184 26.73 22.8%
SupIR 21.78 0.723 0.310 0.198 44.94 27.8% 23.62 0.754 0.308 0.207 27.91 23.6%
DiffTSR 22.10 0.768 0.278 0.168 44.91 44.1% 22.41 0.767 0.300 0.189 31.25 31.7%
DiffBIR 22.44 0.747 0.260 0.201 46.44 37.6% 23.67 0.724 0.262 0.197 23.10 26.6%
OSEDiff 21.86 0.771 0.197 0.127 41.00 26.8% 25.07 0.819 0.201 0.169 20.53 28.9%
DreamClear 22.47 0.772 0.216 0.157 38.97 50.2% 24.10 0.773 0.238 0.191 21.75 21.4%
TSD-SR 21.43 0.754 0.220 0.175 47.21 34.9% 22.79 0.757 0.207 0.194 24.08 26.6%
DiT4SR 20.54 0.764 0.268 0.186 49.79 29.2% 23.16 0.767 0.215 0.159 20.58 23.7%
TADiSR 23.83 0.790 0.286 0.154 44.42 64.7% 24.61 0.796 0.203 0.160 36.61 36.6%
[…]
Table 3: Evaluation of image quality on text regions and text accuracy compared to LR ($\Delta$ OCR-A). TIGER achieves the best performance.
Methods Real-CE UZ-ST
[…]
Real-ESRGAN 21.71 0.824 0.257 0.224 -8.8% 21.25 0.786 0.302 0.266 -4.7%
HAT 23.03 0.862 0.249 0.239 -8.2% 22.18 0.813 0.291 0.276 -3.9%
MARCONet 20.68 0.786 0.267 0.223 -9.8% 17.48 0.690 0.529 0.366 -8.3%
SeeSR 22.74 0.844 0.237 0.212 -27.4% 20.45 0.767 0.297 0.241 -18.9%
SupIR 19.96 0.778 0.326 0.263 -37.0% 20.10 0.756 0.344 0.287 -18.2%
DiffTSR 20.46 0.818 0.275 0.239 -20.7% 17.00 0.665 0.452 0.313 -10.0%
DiffBIR 21.20 0.792 0.278 0.224 -27.2% 20.70 0.757 0.302 0.245 -15.2%
OSEDiff 19.38 0.768 0.269 0.208 -38.0% 21.43 0.798 0.276 0.263 -12.8%
DreamClear 22.74 0.845 0.190 0.168 -14.6% 20.50 0.768 0.317 0.276 -20.3%
TSD-SR 19.33 0.763 0.279 0.234 -29.9% 19.42 0.748 0.296 0.249 -15.1%
DiT4SR 17.95 0.738 0.317 0.244 -35.6% 19.73 0.760 0.273 0.216 -18.0%
TADiSR 23.39 0.855 0.253 0.258 -1.0% 21.59 0.799 0.360 0.336 -5.1%
[…]
Through meticulous annotation and alignment, we obtained 5,036 image pairs with 49,675 text lines. We treat images captured under 200 mm focal lengths as GT, yielding 1,439, 1,798, and 1,799 pairs for the $\times 14.29$, $\times 5.71$, and $\times 2.35$ zoom modes, respectively. Among them, we randomly select 470, 589, and 581 pairs for evaluation under each zoom mode. Each image pair contains one or more text lines. These evaluation sets allow us to assess model performance in more complex and challenging scenarios.
For the UZ-ST-benchmark, we utilize 5 evaluation metrics to examine the quality of image super-resolution on the scale of the full image. Firstly, we adopt Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [wang2019textsr], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], Deep Image Structure and Texture Similarity (DISTS) [ding2020image], and Fréchet Inception Distance (FID) [heusel2017gans] for evaluation of image quality. Specifically, PSNR and SSIM are computed in the pixel space to quantify low-level reconstruction fidelity, LPIPS and DISTS are computed in the feature space to assess perceptual similarity, and FID is employed to evaluate the distributional discrepancy between generated and real images. However, these metrics do not directly reflect the quality or accuracy of super-resolved text. To address this, we crop annotated text regions from the images and compute PSNR, SSIM, LPIPS, and DISTS on these cropped regions, denoted as $\text{PSNR}{cr}$, $\text{SSIM}{cr}$, $\text{LPIPS}{cr}$, and $\text{DISTS}{cr}$. For text accuracy, we apply an OCR model to recognize the text and compare its outputs with the ground-truth annotations using the Levenshtein ratio [yujian2007normalized]:
$\text{OCR-A}=\frac{\text{Len}(s_{pred})+\text{Len}(s_{gt})-\text{Dist}(s_{pred},s_{gt})}{\text{Len}(s_{pred})+\text{Len}(s_{gt})},$ (8)
where $s_{pred}$ stands for the predicted text sequence, $s_{gt}$ denotes annotated ground-truth text sequence, and $\text{Dist}(\cdot,\cdot)$ is the Levenshtein distance between text sequences.
Figure 4: Qualitative Evaluation on Real-CE and UZ-ST.
5 Experiment
5.1 Implementation Details
We combine synthetic data (built upon LSDIR [li2023lsdir] with text rendered via LBTS [tang2023scene] and Real-ESRGAN [wang2021real] degradation) and real data from Real-CE [ma2023benchmark] and UZ-ST. Misaligned pairs in Real-CE are filtered and reannotated, yielding 337 training and 188 testing pairs. Stage 1 uses cropped text regions and is trained using an IDM-based architecture [zhang2024diffusion] with combined segmentation and reconstruction losses. Stage 2 is based on Stable Diffusion 3.5 [esser2024scaling] and employs a tile-based inference strategy [yu2024scaling, hu2025text]; it is pretrained on synthetic data and fine-tuned on real data. For more details, see the Sec. 11.
5.2 Comparison Results
5.2.1 Quantitative Results
We evaluate existing competing methods, including GAN-based image super-resolution approaches such as Real-ESRGAN [wang2021real] and HAT [chen2023activating]; diffusion-based approaches such as SeeSR [wu2024seesr], SupIR [yu2024scaling], DiffBIR [lin2024diffbir], OSEDiff [wu2025one], DreamClear [ai2024dreamclear], TSD-SR [dong2025tsd], and DiT4SR [duan2025dit4sr]; as well as text-focused reconstruction methods such as MARCONet [li2023learning], DiffTSR [zhang2024diffusion], and TADiSR [hu2025text]. Evaluations are performed on the Real-CE Benchmark [ma2023benchmark] and the benchmark described in Sec. 4. To ensure fairness, we fine-tune the released pre-trained models on the training sets of both benchmarks using the official code when available. Following [hu2025text], we integrate the outputs of MARCONet and DiffTSR with HAT-generated results to simulate real-world application scenarios and enable comprehensive full-image evaluation. We evaluate Real-CE at the hardest difficulty level ($\times 4$), while UZ-ST is evaluated across all difficulty levels—with average results reported here. Detailed per-level evaluations appear in the Sec. 14.
As shown in Tab. 2 and Tab. 3 , TIGER outperforms competing methods on both the Real-CE and UZ-ST benchmarks in terms of image quality and text accuracy. TADiSR, limited by the resolution of its cross-attention mechanism, performs poorly under severe degradation and with small text. DiffTSR and MARCONet fail to effectively handle text backgrounds and struggle with text regions that have a large width-to-height aspect ratio. In contrast, our method restores images with high quality and fine-grained text, owing to the decoupling strategy. It achieves OCR-A scores above 0.67 and 0.43, and demonstrates overall superiority in both pixel-level and perceptual accuracy.
Figure 5: Qualitative Results of Ablation Study with stage 2 fixed as the baseline.
5.2.2 Qualitative Results
Fig. 4 shows visual comparisons on Real-CE and UZ-ST. Real-world degradations cause noise, blurred strokes, and distortions in the input images. GAN-based methods reduce noise but fail to fix text structure, while diffusion-based SR methods, though powerful, lack text-structure guidance and often distort strokes, sometimes making text unreadable.
Methods tailored for text image super resolution (e.g., MARCONet and DiffTSR) exhibit limitations in processing long text sequences, often generating distorted or semantically meaningless results. They also lack global semantic guidance, making it hard to blend text regions with backgrounds. As seen in Fig. 4 (a), DiffTSR+HAT (red box) produces text colors inconsistent with the original LR and SR output of HAT. In Fig. 4 (b), it even distorts the face near the word “Ory”. TADiSR works well on lightly degraded images but fails under severe degradation and performs poorly on small text due to cross-attention mask resolution limits. Our method overcomes these issues, restoring high-quality glyph structures even in challenging cases.
5.3 Ablation Study
Table 4: Validation on the effectiveness of UZ-ST. Finetuning improves OCR-A, proving effectiveness.
[…]
OSEDiff w/o UZ-ST 23.50 0.791 0.197 0.162 24.30 22.8%
OSEDiff w/   UZ-ST 25.07 0.819 0.201 0.169 20.53 28.9%
DiT4SR w/o UZ-ST 22.58 0.754 0.252 0.196 26.93 19.3%
DiT4SR w/   UZ-ST 23.17 0.767 0.215 0.159 20.58 23.7%
Ours w/o UZ-ST 22.96 0.782 0.341 0.253 33.78 40.0%
Ours w/   UZST 25.48 0.830 0.196 0.156 20.01 43.0%
Table 5: Ablation on OCR prediction. The model remains strong with null or random OCR text, indicating limited reliance on OCR.
[…]
TADiSR 24.68 0.795 0.362 0.227 52.13 35.5%
Null Text 25.59 0.834 0.188 0.151 28.93 40.4%
Random Text 25.60 0.833 0.188 0.150 28.92 40.3%
Predicted Text (Ours) 25.63 0.834 0.185 0.150 28.82 44.6%
Effectiveness of the training set. Tab. 4 reports consistent OCR-A improvements across different architectures, including the UNet-based OSEDiff and the DiT-based DiT4SR. This demonstrates that even models without explicit mechanisms for text structure modeling can still learn the mapping from low- to high-resolution text using our dataset, further validating its effectiveness. Due to their lack of a dedicated mechanism for text structure modeling, they cannot outperform our methods.
OCR Influence. To investigate how OCR prediction affects the accuracy of our framework, we set the predicted text by OCR in stage 1 as null text or random text and test them on the 35mm subset of UZ-ST. As shown in Tab. 5, even without the OCR model’s predicted text, our method attains a strong OCR-A score of 40.4%, outperforming TAiDSR. This demonstrates that our model in stage 1 benefits from the inherent text structure in the LR input to restore text structure and is not overly dependent on the OCR model.
Effectiveness of the TIGER component. We set stage 2 as the fixed baseline for image restoration and validate the effectiveness of our framework through 4 ablated variants:
  • Empty Mask Guidance. As shown in Tab. 4, empty masks provide zero structure guidance, and the model can only rely on the LR for text restoration. Therefore, the OCR-A drops from 67.3% to 59.5%.
  • Standard Font Guidance. As shown in the second column of Fig. 5, the standard font, while structurally correct, provides weak guidance due to style and position mismatches. As a result, the restored text exhibits poor accuracy compared to TIGER, with the OCR-A dropping from 67.1% to 55.3%, as shown in Tab. 6.
  • SAM-TS Extraction [hu2025text, ye2024hi]. While this improves visual quality by better aligning output masks with the LR image, SAM-TS can only extract distorted structures and can’t recover degraded text (Fig. 5, third column), leading to suboptimal performance compared with TIGER, with OCR-A dropping from 0.671 to 0.579.
  • LDM Guidance. This variant uses a latent diffusion model conditioned on LR to restore text structure, achieving better accuracy by reconstructing or compensating for lost text. However, without separate learning from synthetic and real-world data, it struggles with high-quality text structures in real-world scenarios. As shown in Fig. 5, fourth column, the masks remain noisy under real-world degradations, leading to lower accuracy than TIGER, with OCR-A dropping from 67.1% to 60.1% as reported in Tab. 6.
Our method uses RGB output to jointly learn from real-world degradations and accurate synthetic masks, yielding high-quality masks and achieving the highest accuracy.
Table 6: Ablation study on Real-CE with stage 2 fixed as the baseline. From top to bottom, we compare the performance of using text masks rendered with a standard font, extracted using SAM-TS, reconstructed with latent diffusion model conditioned on LR, and reconstructed with our text restoration pipeline. Our pipeline faithfully restores glyph structures, yielding the highest accuracy.
[…]
Empty Mask 23.33 0.792 0.225 0.172 42.93 59.5%
Std Font Guidance 23.31 0.786 0.249 0.164 41.83 55.3%
SAM-TS Extraction 23.91 0.835 0.212 0.147 41.22 57.9%
LDM Guidance 22.42 0.798 0.222 0.147 43.08 60.1%
[…]
6 Conclusion
In this paper, we delve into the extensively researched problem of scene text super-resolution. To address this challenge, we propose a novel approach, TIGER, that decouples glyph structure restoration from image enhancement. For the restoration of text structure, we propose the text restoration pipeline, which enables a 2-phase training strategy to fully take advantage of both synthetic and real-world data. For the image enhancement, we propose the image enhancement pipeline, which effectively utilizes the glyph structure to restore both text fidelity and image details coherently in full-image super-resolution. In terms of training data and benchmark, we present the UltraZoom-ST, the first real-world scene text dataset that contains extreme zooming mode, offering more challenging scenarios for the field. Extensive experiments on Real-CE and UZ-ST demonstrate the superiority of TIGER over existing methods.
[…]
Supplementary Material
Table 7: Statistics of dataset size and line count in subsets of UltraZoom-ST.
Subset image count line count mean lines/img img >> 5 lines
14 mm 1,439 15,073 10.47 754
35 mm 1,798 17,263 9.60 867
85 mm 1,799 17,339 9.64 869
Table 8: Alignment comparison.
Alignment Method PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ NCC $\uparrow$ AKD $\downarrow$
RealSR 13.87 4002.65 0.5072 0.3285 546.78
SIFT Alignment (Single time) 12.77 6332.79 0.4896 0.4313 546.78
[…]
Table 9: Efficiency analysis.
Methods Flops (GFLOPs) Speed (ms) OCR-A
HAT 6670.32 1086.68 37.9%
DiffTSR 58502.14 8610.59 31.7%
DiT4SR w/o llava 160787.14 17385.14 23.7%
DreamClear w/o llava 412843.67 83193.81 21.4%
TADiSR 4497.96 342.32 36.6%
TIGER (ours) Stage 1 6215.85 663.5 43.0%
TIGER (ours) Stage 2 3734.01 360.06
Table 10: Ablation on the effect of stochastic OCR outputs on the performance of our method. The performance gains as the randomness of the OCR output drops. Even under 100% random OCR output, our method still outperforms TADiSR, proving low reliance on OCR.
[…]
TADiSR 24.68 0.795 0.362 0.227 52.13 35.5%
TIGER (Ours) performance under different randomness of OCR output.
100% Random Text 25.60 0.833 0.188 0.150 28.92 40.3%
50% Random Text 25.60 0.833 0.189 0.151 28.92 40.6%
20% Random Text 25.63 0.834 0.185 0.150 29.3 43.7%
0% Random Text (Ours) 25.63 0.834 0.185 0.150 28.82 44.6%
Table 11: Evaluation Results on Both Image Quality and Text Accuracy on UltraZoom-ST.
[…]
85mm ($\times 2.35$) Real-ESRGAN 25.84 0.854 0.139 0.130 28.25 23.59 0.843 0.170 0.168 60.8%
HAT 27.33 0.873 0.128 0.133 27.21 24.76 0.870 0.139 0.146 62.0%
MACRONet 22.90 0.804 0.234 0.173 40.19 17.92 0.693 0.515 0.360 54.2%
SeeSR 25.02 0.826 0.167 0.159 32.31 21.62 0.796 0.245 0.207 28.2%
SupIR 25.58 0.823 0.202 0.160 29.33 21.91 0.808 0.239 0.214 33.3%
DiffTSR 23.48 0.805 0.230 0.157 35.77 17.92 0.688 0.392 0.285 48.2%
DiffBIR 25.25 0.793 0.175 0.152 24.57 22.26 0.803 0.218 0.190 36.8%
OSEDiff 26.38 0.852 0.153 0.137 23.92 22.84 0.833 0.200 0.204 41.0%
DreamClear 25.31 0.811 0.169 0.151 24.19 21.44 0.790 0.253 0.221 26.8%
TSD-SR 23.99 0.800 0.161 0.165 30.40 20.62 0.786 0.232 0.218 37.8%
DiT4SR 24.71 0.810 0.173 0.143 29.15 21.25 0.800 0.221 0.193 48.2%
TADiSR 26.31 0.858 0.177 0.149 30.47 23.55 0.856 0.175 0.178 55.7%
Ours 27.00 0.871 0.123 0.120 22.34 24.31 0.860 0.135 0.142 63.2%
35mm ($\times 5.71$) Real-ESRGAN 23.70 0.785 0.218 0.184 42.62 20.86 0.782 0.272 0.235 34.1%
HAT 25.05 0.819 0.207 0.182 41.32 21.98 0.816 0.250 0.241 34.6%
MACRONet 22.28 0.774 0.284 0.199 49.14 17.43 0.694 0.523 0.361 31.5%
SeeSR 23.76 0.795 0.198 0.175 37.92 20.55 0.778 0.271 0.222 25.2%
SupIR 23.49 0.746 0.302 0.204 41.80 20.00 0.757 0.335 0.282 24.0%
DiffTSR 22.57 0.773 0.281 0.184 45.78 16.99 0.671 0.443 0.312 31.2%
DiffBIR 23.76 0.731 0.245 0.196 35.99 20.82 0.770 0.275 0.234 28.6%
OSEDiff 25.35 0.827 0.192 0.166 29.06 21.62 0.809 0.263 0.263 29.9%
DreamClear 24.20 0.778 0.209 0.177 32.08 20.66 0.779 0.283 0.251 23.9%
TSD-SR 22.93 0.757 0.200 0.199 37.61 19.64 0.760 0.282 0.246 27.3%
DiT4SR 23.45 0.774 0.204 0.158 31.73 20.12 0.779 0.250 0.208 26.5%
TADiSR 24.68 0.795 0.362 0.227 52.13 21.57 0.798 0.366 0.347 35.5%
Ours 25.63 0.834 0.185 0.150 28.82 22.14 0.823 0.204 0.194 44.6%
14mm ($\times 14.29$) Real-ESRGAN 22.09 0.718 0.423 0.284 89.22 18.87 0.721 0.503 0.427 11.5%
HAT 22.66 0.738 0.451 0.299 89.58 19.27 0.738 0.533 0.484 12.2%
MACRONet 21.00 0.716 0.422 0.254 79.67 17.01 0.680 0.553 0.382 10.1%
SeeSR 21.79 0.731 0.310 0.228 64.31 18.87 0.717 0.396 0.308 13.2%
SupIR 21.38 0.681 0.447 0.268 70.09 17.99 0.690 0.487 0.383 11.0%
DiffTSR 20.88 0.713 0.412 0.233 74.07 15.87 0.630 0.538 0.348 12.0%
DiffBIR 21.61 0.630 0.392 0.255 69.64 18.61 0.685 0.438 0.328 11.3%
OSEDiff 23.11 0.768 0.274 0.214 49.76 19.44 0.740 0.385 0.334 12.7%
DreamClear 22.45 0.719 0.364 0.260 62.95 19.14 0.725 0.440 0.377 11.7%
TSD-SR 21.14 0.705 0.273 0.225 57.07 17.67 0.686 0.393 0.293 11.9%
DiT4SR 20.88 0.707 0.282 0.181 47.45 17.36 0.687 0.367 0.253 11.7%
TADiSR 22.43 0.721 0.555 0.320 98.44 19.20 0.730 0.585 0.518 14.4%
Ours 23.41 0.774 0.301 0.209 54.74 19.73 0.748 0.374 0.322 16.0%
Total RealEsrGAN 23.99 0.790 0.248 0.194 30.60 21.25 0.786 0.302 0.266 37.1%
HAT 25.17 0.815 0.249 0.198 30.12 22.18 0.813 0.291 0.276 37.9%
MARCONet 22.13 0.768 0.306 0.205 34.28 17.48 0.690 0.529 0.366 33.4%
SeeSR 23.64 0.788 0.219 0.184 26.73 20.45 0.767 0.297 0.241 22.8%
SupIR 23.62 0.754 0.308 0.207 27.91 20.10 0.756 0.344 0.287 23.6%
DiffTSR 22.41 0.767 0.300 0.189 31.25 17.00 0.665 0.452 0.313 31.7%
DiffBIR 23.67 0.724 0.262 0.197 23.10 20.70 0.757 0.302 0.245 26.6%
OSEDiff 25.07 0.819 0.201 0.169 20.53 21.43 0.798 0.276 0.263 28.9%
DreamClear 24.10 0.773 0.238 0.191 21.75 20.50 0.768 0.317 0.276 21.4%
TSD-SR 22.79 0.757 0.207 0.194 24.08 19.42 0.748 0.296 0.249 26.6%
DiT4SR 23.16 0.767 0.215 0.159 20.58 19.73 0.760 0.273 0.216 23.7%
TADiSR 24.61 0.796 0.203 0.160 36.61 21.59 0.799 0.360 0.336 36.6%
Ours 25.48 0.830 0.196 0.156 20.01 22.22 0.814 0.228 0.212 43.0%
7 Reproducibility Statement
To ensure reproducibility, we have made the following efforts:
  1. 1.
    We will release our code and dataset.
  2. 2.
    We provide implementation details in Sec. 5.1 and Sec. 11, including the training process and selection of hyper-parameters.
  3. 3.
    We provide details on evaluation metrics and dataset preparation in Sec. 4 and Sec. 10, and the code and data will be made available along with it.
8 Ethics Statement
This work focuses on improving scene text image super-resolution to support beneficial applications such as enhancing accessibility, document restoration, and navigation assistance. However, we acknowledge potential risks, including misuse in privacy-sensitive contexts (e.g., recovering text from personal or social media images) or unintended deployment in surveillance. To mitigate such risks, users are encouraged to combine our methods with privacy-preserving techniques such as watermarking or selective inpainting. Our UZ-ST dataset was collected from public, non-sensitive scenes under ethical guidelines, without including personally identifiable or private information. We believe the societal benefits of improved text image restoration outweigh potential risks, provided that the technology is applied responsibly.
9 LLM Acknowledgments
We thank ChatGPT (GPT-5) and DeepSeek for the help in refining the language and improving clarity. All wording and factual content were reviewed and approved by the authors.
10 Dataset Collection
We use VIVO X200 Ultra to collect images for 4 separate focal lengths (14 mm, 35 mm, 84 mm, and 200 mm). We first use PP-OCRV5 [cui2025paddleocr30technicalreport] for the rough annotation, then we manually filter the images and annotations. During the filtering and annotation, each image undergoes the following rules:
  • Width or height of the image should be no less than 256.
  • Height of the text should not be less than 32 pixels.
  • Score of OCR recognition of the text should not be lower than 0.9.
  • Content of the text should not be empty or consist solely of whitespace.
11 More Implementation Details
Dataset Settings. For training, we combine a synthetic dataset with real paired datasets (Real-CE [ma2023benchmark] and UZ-ST). Our synthetic dataset builds upon LSDIR [li2023lsdir], containing 27,000 triplets $(x_{H},x_{L},x_{m})$. We render text on the GT of LSDIR and the corresponding text mask using the LBTS [tang2023scene], then apply Real-ESRGAN degradation wang2021real to generate LR images. Since there are misaligned image pairs in Real-CE [hu2025text, zhang2024diffusion] and text lines not annotated, we filter out misaligned image pairs and reannotate the images manually. In the end, we obtain 337 training pairs and 188 testing pairs from the Real-CE dataset. Images under 13mm and 52mm focal lengths are considered as LR $x_{L}$ and GT $x_{H}$, respectively. Following hu2025text, we use SAM-TS [ye2024hi] to obtain the text mask $x_{m}$ from $x_{H}$. UZ-ST triplets $(x_{H},x_{L},x_{m})$ are processed identically. For stage 1 training, we use cropped text regions from the dataset, while for stage 2, we use full images. Following [DBLP:conf/iclr/TuoXHGX24], we use PP-OCRv3 [DBLP:journals/corr/abs-2206-03001] to extract strings from the text region of predicted SR for evaluation of OCR-A.
Training Details. For stage 1, we build our model based on the IDM baseline of DiffTSR [zhang2024diffusion]. We set $\lambda_{td}$, $\lambda_{Seg}$, $\lambda_{Focal}$, and $\lambda_{Dice}$ as $1$, $0.1$, 2020, and $1$ respectively. In stage 2, our model is built upon Stable Diffusion 3.5 medium [esser2024scaling] and follows a similar tile-based inference strategy to TADiSR [hu2025text] and SUPIR [yu2024scaling]. The timestep $t$ is set to $150$. The conditioning scale of the ControlNet is set to 1.0. $\lambda_{l2}$, $\lambda_{LPIPS}$, and $\lambda_{edge}$ are set to $1$, 55, 100100 when using the synthetic dataset for pretraining. Then, $\lambda_{edge}$ is set to 0 and the remaining coefficients are kept unchanged when training on real paired datasets. We train both stages of the model using the AdamW [loshchilov2017decoupled] optimizer and set the learning rate to $5\times 10^{-5}$ and $5\times 10^{-6}$ for stages 1 and 2 separately. All experiments are conducted on NVIDIA H20 GPUs. For stage 1, we train the model on synthetic and real datasets for 8 epochs, then finetune it on synthetic data for 2 epochs. For stage 2, we first pretrain the model on the synthetic dataset for 50 epochs. Then we train on the real paired datasets for 50 epochs.
12 More details of UZ-ST
In Tab. 7, we provide detailed statistics on the composition of the UZ-ST dataset. Additionally, in Fig. 10, we present some example images from the dataset.
Figure 6: Alignment method comparison.
13 Alignment Comparison
We compare our proposed Cascade Coarse-to-Fine alignment strategy with the alignment method proposed by RealSR [cai2019toward] and single-time alignment using SIFT. During the comparison, we implement our strategy using SIFT and utilize the raw images from the 35mm dataset of our proposed UZ-ST dataset for evaluation. For metrics, we adopt PSNR, Mean Squared Error (MSE), SSIM [wang2019textsr], Normalized Cross-Correlation (NCC), and Average Keypoint Distance (AKD) to assess the alignment quality of the aligned image.
As shown in Tab. 8 and Fig. 6, pixel-based optimization methods break when handling heavily degraded images due to significant loss of details. Keypoint-based alignment like SIFT suffers from the lack of valid keypoint for matching. While our alignment method outperforms all other methods and achieves the best results, this is due to the progressive cascade alignment strategy, which breaks the hard alignment process into simpler tasks, facilitating the bridging between images taken under different focal lengths.
14 Additional results
We provide detailed quantitative results of UZ-ST in Tab. 11. Additional qualitative results are provided in Fig. 7 and Fig. 8. Fig. 9 also shows the ability of our method in handling other languages like Japanese and Korean.
We also conduct additional ablation experiment on how the randomness of OCR output affect our method. As shown in Tab. 10, the text accuracy gains as the randomness of the OCR output declines. However, even under complete random OCR prediction, our method still manage to outperform TADiSR thanks to the LR condition still guiding the stage 1 to generate text structures. This proves that the reliance of our method on the OCR model is limited.
15 Efficiency Analysis
As shown in Tab. 9, stage 1 is a standard diffusion process that takes multiple steps in inference, the efficiency of our model may be suboptimal compared to one-step methods. However, our method achieves state-of-the-art performance in OCR-A, which cannot be easily obtained by extending inference time.
Figure 7: Additional qualitative results of TIGER.
Figure 8: Additional qualitative results of TIGER.
Figure 9: Demonstration of our method under different languages. (Zoom in for more details.)
Figure 10: Detailed examples of UZ-ST.